# ORBIT: An Object Property Reasoning Benchmark for Visual Inference Tasks

### v1

@@ Conclusions
This paper developed an evaluation framework for evaluating object-centric reasoning in images. The framework considers three distinct image types (photographic, animated, and AI-generated), three reasoning levels (direct recognition, property inference, counterfactual), and four property types (physical, taxonomic, functional, relational). We instantiated this framework into ORBIT, a VQA benchmark with over 1,000 questions for 360 images, split uniformly across the evaluation dimensions. Experiments with 12 state-of-the-art VLMs revealed low accuracy, with most models performing far below humans, and the best only achieving 40% accuracy. VLMs struggle particularly with realistic (photographic) images, counterfactual reasoning about physical and functional properties, and infrequent counts over 5. Humans also struggle with photographic images and physical properties, but excel at counterfactual questions and have no frequency effects on counts.
@@
Three key limitations should be addressed in future work. First, as generative AI still suffers from limited accuracy and diversity in autonomous question generation, which makes scaling up object abstraction benchmarks require substantial human effort. Novel approaches, such as metamorphic testing (Yuan et al. 2021), are necessary to grow ORBIT systematically in size and coverage of additional object properties, question types beyond counting, and open-ended answers. Second, since annotators are unrestricted in their question design, they may introduce linguistic ambiguity, leading to disagreements across annotators (e.g., whether zebras are considered black animals). While we addressed ambiguity in an explicit quality assurance step, during which sources of disagreements are analyzed and converted into more precise guidelines, some disagreements remain. Future work should develop generalizable, well-understood guidelines for object abstraction and reasoning. Finally, while ORBIT points to fundamental limitations of 12 popular and diverse VLMs, the set of baselines is inherently incomplete. Future work should explore additional specialized reasoners beyond SpatialVLM, e.g., mixtures of experts (Lin et al. 2024) and derivation of object schemas (Hsu et al. 2024).
Acknowledgements
This paper is based on Abhishek Kolari’s MSc AI project at Vrije Universiteit Amsterdam with the same title. MK and FI were funded by the AiNed project Human-Centric AI Agents with Common Sense supported by the Dutch Organisation for Scientific Research (NWO), under grant no. NGF.1607.22.044. FdH was funded by the Hybrid Intelligence Centre, a 10-year programme supported by the Dutch Ministry of Education, Culture and Science through NWO (grant no. 024.004.022). The evaluations were performed on the DAS-6 cluster (Bal et al. 2016).
Appendix A Appendix
Appendix B Benchmark Examples
Image Types
Figure 5 shows additional examples from ORBIT for three image types: Photographic, Animated, and AI-generated.
Figure 5: Examples from ORBIT, for three image types: Photographic (top-left), Animated (top-right), and AI-generated (bottom). Text highlighted in blue is information stored in the benchmark and not passed to the model during evaluation. GT stands for ground truth, RC is reasoning complexity, and OP for object properties.
Object Property Dimensions
Reasoning Complexity
Table 5 shows examples of the questions asked at the different reasoning levels. Direct recognition has questions from the taxonomical and physical object property dimensions; property inference has questions from the functional and relational dimensions; and counterfactual reasoning has questions from any of the four property dimensions.
Appendix C Prompt Strategies
Question Generation
Output Formatting
@@ Appendix C Prompt Strategies
Table 7: The 15 tags (i.e., locations and activities) taken from the MacGyver dataset (Tian et al. 2025), extended with 11 of our own, for a total of 26 unique image themes. They are broadly divided into Indoors/Household, Neutral, and Outdoors.
AI-Generated images
@@
Table 8 shows the prompt examples used while generating AI-based images for different image themes.
Image Theme Prompt Example Campsite settings (Outdoors) Create image Generate a photorealistic dusk campsite scene with 5 tents, a few of them illuminated from within, a controlled campfire with 3 people sitting on logs around it and a dog sleeping nearby, hiking backpacks leaning against a nearby tree, a lantern hanging from a branch, and stars beginning to appear in the darkening sky. Include a mountain silhouette in the background. Strictly keep the count of total objects in the image to a maximum of 10. Gym (Indoors/Household) Create image Generate a photorealistic image of a gym scene with sports equipment. Let the gym contain 3 treadmills, 2 elliptical machines, a few kettlebells on the floor, 2 resistance bands hanging on the wall, and a bench press rack. Let it be a realistic setting with 2 humans working out, one person on the treadmill and another person on the elliptical machine. The gym can have a clock and TV hanging on the wall somewhere showing a workout video. Include large windows with natural light and a small plant in the corner. Strictly keep count of total objects in the image to a maximum of 10. Beach cleanup (Outdoors) Generate a photorealistic image of a beach cleanup scene with 4 volunteers in bright t-shirts collecting trash at sunset in the backdrop. Let the trash include these objects: 5 glass bottles, 6 trash bags (one in each hand of a volunteer and 2 placed near the ground), a collection bin, reusable gloves, and garbage pickers. Let it be a realistic setting, having the ocean with gentle waves in the background. Strictly keep the count of total objects in the image to a maximum of 10. Garage (Neutral) Create image Create a photorealistic image of a vehicle garage scene during a vehicle maintenance session. Show 3 cars, one car on a hydraulic lift with its hood open, 2-3 essential tools spread on a workbench, an oil change in progress, spare parts in organised containers, and a service manual. Let the setting be realistic with ceiling lights for the garage, let the car being serviced be in good condition, and let the other two cars in the background be scrap and old, which need to be rebuilt with loose tires leaning next to them. Let one person work on the car that is being serviced and the other person in the backdrop next to the other 2 cars. Strictly keep count of total objects in the image to a maximum of 10. Kitchen (Indoors/Household) Create image Generate an image of a kitchen countertop during the preparation of a complex dish. Show fresh ingredients arranged in 3 small ceramic bowls, a wooden cutting board with chopped herbs, a professional knife set, a simmering pot on the stove, a recipe book, a glass filled with water, and 2 people collaborating on the cooking. Include details like steam, a kitchen sink next to the stove, and ambient kitchen lighting. Let it be a realistic setting. Strictly keep the count of total objects in the image to a maximum of 10.
Table 8: Prompt examples used to generate AI images across different themes, along with corresponding tags (shown in brackets) used in constructing the MacGyver dataset. Prompts with the Create Image tool are generated using GPT-4o, while those without it are created with Grok 3.
Appendix D Ground Truth Distribution
@@
Appendix E Image Theme Tags
During image collection (Figure 2), to direct our search we made use of image theme tags inspired by the tags used in curating the Macgyver dataset (Tian et al. 2025). Table 7 shows the list of 26 unique image theme tags used for the curation of ORBIT. The process of selecting a tag was influenced by the chance of a particular theme having more variety of objects. The tags used for the MacGyver dataset were plentiful out of which 15 tags were in common, namely: bedroom, kitchen, library, meeting room, gym, wardrobe, garage, classroom, beach cleanup, camping, construction, gardening, zoo, park, and farm. The remaining 9 tags came into fruition off our own efforts.
Appendix F Inductive Bias
The BLIP2 model series consistently defaults to specific output values, showing possibilities of inductive bias present in the models. The OPT 2.7B variant goes for the count 1, the OPT 6.7B variant produces either 1 or 3 as the output, and the FLan T5 xxL gives an output as 0 or no value or an answer not abiding by the required format (numeric value), resulting in a final output as 0 set according to our answer extraction method. The frequency distribution of these outputs is shown in Figure 7. We tried adjusting prompts and decoding parameters, such as the number of maximum new tokens, but none helped mitigate the issue, highlighting the potential inductive biases present in models (Wang and Wu 2023; Kraaijveld et al. 2025).
Appendix G Cumulative Accuracies with Tolerance
Appendix H Visual Depiction of Errors
ORBIT requires models to have good low-level perception and the ability to perform fine-grained analysis. A case study on the errors models typically make is visually depicted using an image of photographic type. Figure 9 shows an instance from ORBIT, with the ground truth values provided after the question in the yellow cell and the depiction of object counts through green bounding boxes. The bounding boxes are manually annotated on the images using Canva. [^1] https://www.canva.com/ To showcase the output of the model, we consider the open-source model, Gemma 3 (27B), for this study. Figure 10 displays the model’s predicted values and the probable object selection, with green (correct) and red (incorrect) bounding boxes based on detailed reasoning explicitly requested for this image.
@@ Appendix H Visual Depiction of Errors
The model fails to detect objects that are partially hidden, indicating weak low-level perception. It misinterprets perspective and geometric layout and struggles to distinguish objects that are closely clustered together. Despite the incorrect answers, the model’s internal reasoning is logically consistent. This study highlights the importance of evaluating not only the final count, but also the quality of visual grounding and traceability of reasoning for tasks like those in ORBIT, providing potential for future work.
@@
Appendix I Human Evaluation
An analysis of human answers indicates that the sources of incorrect responses or disagreements between annotators could be one of the following: 1) Ambiguity in image context: Some mistakes are because of partially visible objects, as people were unsure whether to count them. 2) Uncertainty in definitions: Some mistakes appear from uncertainty in the definition of objects. For instance, it was unclear to human evaluators whether lamps should be considered electronic devices, what qualifies as a metal object, or which items are included when asked about objects with arms. While we agreed on these guidelines between the original annotators, this was not used to instruct the human evaluators. 3) Lack of attention: We also observed some mistakes in relatively straightforward questions, which we attribute to humans not examining the image carefully.
This analysis suggests that human performance could be improved in the future by providing clear guidance for questions on the first and second issues. As a result, we would expect higher accuracy and agreement among human annotators for the first two error categories.
@@ Appendix I Human Evaluation
Appendix J VG-based Dataset Curation
Figure 11 shows two examples of question-answer pairs generated automatically from Visual Genome.

### v2

@@ 7 Conclusions
This paper developed a commonsense evaluation framework for evaluating object property reasoning in images. The framework considered three distinct image types (photographic, animated, and AI-generated), three reasoning levels (direct recognition, property inference, counterfactual), and four property types (physical, taxonomic, functional, relational). We instantiated this framework on two VQA benchmarks: OPTICS-CNT, with over 1,000 counting questions across 360 images, and OPTICS-CMP, with over 2,000 comparison questions across as many images. Experiments with 12 state-of-the-art VLMs revealed low counting accuracy, with most models performing far below human performance and the best achieving only 40% accuracy. VLMs struggle particularly with realistic (photographic) images, counterfactual reasoning about physical and functional properties, and infrequent counts over five. Humans also struggle with photographic images and physical properties but excel at counterfactual questions and have no frequency effects on counts. Interestingly, when asked to compare images based on their counts, most VLMs performed better for higher differences, but their performance on very large differences remained relatively modest. Analyzing closely the errors made by humans and VLMs revealed that, while both make a sizeable proportion of perceptual mistakes, VLMs are dominated by unnatural errors, whereas humans struggle with semantic ambiguity. This indicates that VLMs experience limitations on a combination of perception and reasoning abilities and unexpected hallucinations.
@@
Future work should address three emerging limitations. First, generative AI still suffers from limited accuracy and diversity in autonomous question generation, causing scaling up object reasoning benchmarks to require substantial human effort. Our attempt to generate a scalable version based on richly annotated data (§B.5) showed that automatically creating such benchmarks remains challenging. Novel approaches, such as metamorphic testing Yuan et al. (2021), are necessary to grow OPTICS-CNT systematically in size and coverage of additional object properties, question types beyond counting and comparison, and open-ended answers. Second, since annotators are free to design their own questions, they may introduce semantic ambiguity, leading to disagreements (e.g., whether zebras are considered black animals). Although we addressed ambiguity through an explicit quality assurance step that analyzes sources of disagreement and translates them into more precise guidelines, some disagreements remain. Future work should develop generalizable and well-understood guidelines for object abstraction and reasoning. Finally, while OPTICS-CNT and OPTICS-CMP point to fundamental limitations of popular and diverse VLMs, the set of baselines is inherently incomplete. Future work should explore additional specialized reasoners beyond SpatialVLM, e.g., a referring expression counting method to address compositional attribute filters and counting Dai et al. (2024), mixtures of experts (Lin et al., 2024), derivation of object schemas Hsu et al. (2024), and methods for populating situational scene graphs following predefined slots Sugandhika et al. (2025).
Appendix A Appendix: Technical Details
@@
Image Domain Prompt Example Campsite settings (Outdoors) Create image Generate a photorealistic dusk campsite scene with 5 tents, a few of them illuminated from within, a controlled campfire with 3 people sitting on logs around it and a dog sleeping nearby, hiking backpacks leaning against a nearby tree, a lantern hanging from a branch, and stars beginning to appear in the darkening sky. Include a mountain silhouette in the background. Strictly keep the count of total objects in the image to a maximum of 10. Gym (Indoors/Household) Create image Generate a photorealistic image of a gym scene with sports equipment. Let the gym contain 3 treadmills, 2 elliptical machines, a few kettlebells on the floor, 2 resistance bands hanging on the wall, and a bench press rack. Let it be a realistic setting with 2 humans working out, one person on the treadmill and another person on the elliptical machine. The gym can have a clock and TV hanging on the wall somewhere showing a workout video. Include large windows with natural light and a small plant in the corner. Strictly keep count of total objects in the image to a maximum of 10. Beach cleanup (Outdoors) Generate a photorealistic image of a beach cleanup scene with 4 volunteers in bright t-shirts collecting trash at sunset in the backdrop. Let the trash include these objects: 5 glass bottles, 6 trash bags (one in each hand of a volunteer and 2 placed near the ground), a collection bin, reusable gloves, and garbage pickers. Let it be a realistic setting, having the ocean with gentle waves in the background. Strictly keep the count of total objects in the image to a maximum of 10. Garage (Neutral) Create image Create a photorealistic image of a vehicle garage scene during a vehicle maintenance session. Show 3 cars, one car on a hydraulic lift with its hood open, 2-3 essential tools spread on a workbench, an oil change in progress, spare parts in organised containers, and a service manual. Let the setting be realistic with ceiling lights for the garage, let the car being serviced be in good condition, and let the other two cars in the background be scrap and old, which need to be rebuilt with loose tires leaning next to them. Let one person work on the car that is being serviced and the other person in the backdrop next to the other 2 cars. Strictly keep count of total objects in the image to a maximum of 10. Kitchen (Indoors/Household) Create image Generate an image of a kitchen countertop during the preparation of a complex dish. Show fresh ingredients arranged in 3 small ceramic bowls, a wooden cutting board with chopped herbs, a professional knife set, a simmering pot on the stove, a recipe book, a glass filled with water, and 2 people collaborating on the cooking. Include details like steam, a kitchen sink next to the stove, and ambient kitchen lighting. Let it be a realistic setting. Strictly keep the count of total objects in the image to a maximum of 10.
Table 5: Prompt examples used to generate AI images across different domains, along with corresponding tags (shown in brackets) used in constructing the MacGyver dataset. Prompts with the Create Image tool are generated using GPT-4o, while those without it are created with Grok 3.
@@
Table 6: The 15 domains (i.e., locations and activities) taken from the MacGyver dataset (Tian et al., 2025), extended with 11 of our own, for a total of 26 unique image domains. They are broadly divided into Indoors/Household, Neutral, and Outdoors.
A.1 Prompt Strategies
A.1.1 Question Generation
A.1.2 Output Formatting
A.1.3 AI-Generated images
@@
Table 5 shows the prompt examples used while generating AI-based images for different image domains.
A.2 Image Domains
During image collection (Figure 2), to direct our search, we made use of image domains inspired by the tags used in curating the MacGyver dataset Tian et al. (2025). The three broad categories in this dataset are: indoors/household (e.g., kitchen), neutral (e.g., garage), and outdoors (e.g., zoo). Table 6 shows the list of 26 unique image domains used for the curation of OPTICS. The process of selecting a domain was influenced by the chance of a particular domain having a greater expected variety of objects. The domain tags used for the MacGyver dataset were plentiful, out of which 15 were in common, namely: bedroom, kitchen, library, meeting room, gym, wardrobe, garage, classroom, beach cleanup, camping, construction, gardening, zoo, park, and farm. The remaining 11 tags resulted from our exploration.
A.3 Benchmark Examples
A.3.1 Image Types
Figure 8 shows additional examples from OPTICS-CNT for three image types: Photographic, Animated, and AI-generated.
Figure 8: Examples from OPTICS-CNT, for three image types: Photographic (top), Animated (middle), and AI-generated (bottom). Text highlighted in blue is information stored in the benchmark and not passed to the model during evaluation. GT stands for ground truth, RC for reasoning complexity, and OP for object properties.
A.3.2 Object Property Dimensions
A.3.3 Reasoning Complexity
Table 8 shows examples of the questions asked at the different reasoning levels. Direct recognition has questions from the physical and taxonomical object property dimensions; property inference has questions from the functional and relational dimensions; and counterfactual reasoning has questions from any of the four property dimensions.
A.3.4 OPTICS-CMP
Figure 9 shows an example from OPTICS-CMP where we merge two images from OPTICS-CNT into a single image. Figure 10 visualizes the sample distribution across count-difference levels and question categories.
Figure 9: Example from OPTICS-CMP with a count difference of 2.
Figure 10: Sample Distribution of OPTICS-CMP.
A.4 Ground Truth Distribution
@@
A.5 Examples of Model Errors
OPTICS requires models to have good low-level perception and the ability to perform fine-grained analysis. A case study on the errors models typically make is visually depicted using an image of photographic type. Figure 12 shows an instance from OPTICS-CNT, with the ground truth values provided after the question in the yellow cell and the depiction of object counts through green bounding boxes. The bounding boxes are manually annotated on the images using Canva. [^2] https://www.canva.com/ To showcase the output of the model, we consider the open-source model, Gemma 3 (27B), for this study. Figure 13 displays the model’s predicted values and the probable object selection, with green (correct) and red (incorrect) bounding boxes based on detailed reasoning explicitly requested for this image.
@@ A.5 Examples of Model Errors
The model fails to detect objects that are partially hidden, indicating weak low-level perception. It misinterprets perspective and geometric layout and struggles to distinguish objects that are closely clustered together. Despite the incorrect answers, the model’s internal reasoning is logically consistent. This study highlights the importance of evaluating not only the final count, but also the quality of visual grounding and traceability of reasoning for tasks like those in OPTICS, providing potential for future work.
@@
Appendix B Appendix: Complementary Results
B.1 Inductive Bias
The BLIP2 model series consistently defaults to specific output values, suggesting the presence of inductive bias in the models. The OPT 2.7B variant goes for the count 1, the OPT 6.7B variant produces either 1 or 3 as the output, and the Flan T5 xxL gives an output of 0 or no value or an answer not abiding by the required format (numeric value), resulting in a final output of 0 set according to our answer extraction method. The frequency distribution of these outputs is shown in Figure 14. We tried adjusting prompts and decoding parameters, such as the number of maximum new tokens, but none helped mitigate the issue, highlighting the potential inductive biases present in models (Wang and Wu, 2023; Kraaijveld et al., 2025).
B.2 Cumulative Accuracies with Tolerance
B.3 OPTICS-CMP results
Table 9 shows the results for seven instruction-tuned models evaluated on OPTICS-CMP.
Model Micro
Acc
($\uparrow$)
Macro
Acc
($\uparrow$)
Qwen2.5 (3B) 55.14 56.30
Qwen2.5 (7B) 57.54 58.10
Qwen2.5 (32B) 68.33 72.39
InternVL3 (8B) 62.02 64.35
InternVL3 (14B) 46.84 64.19
Gemma3 (27B) 54.52 64.48
GPT-4o mini 60.34 64.45
Model average 57.81 63.46
Table 9: Zero-shot results on OPTICS-CMP. The micro accuracy and macro accuracy (average across count differences (1-9)) are reported. The highest and second-highest model results are highlighted in bold and underlined, respectively. ($\uparrow$) indicates that higher values are better.
B.4 Error Analysis of Humans and Qwen
The raw results of this analysis are provided in Table 10.
Semantic
ambiguity
Visual
ambiguity
Perception
errors
Unnatural
errors
Human 16 3 9 0
Qwen2.5 (32B) 8 1 8 9
Table 10: Results from analyzing the errors made by humans and the model.
@@ B.4 Error Analysis of Humans and Qwen
B.5 Visual Genome-based Dataset Curation
Figure 16 shows two examples of question-answer pairs generated automatically from Visual Genome. We pose the following question: Can densely annotated datasets be used to scale up OPTICS-CNT automatically? Visual Genome (VG) Krishna et al. (2017) is a popular dataset with high-quality structured connections between image regions, language descriptions, and WordNet Miller (1995) synsets. To investigate whether the annotations of VG can support the generation of large-scale instantiations of OPTICS, we devised a procedure that generates taxonomic questions by manipulating the WordNet synsets. This experiment exposed three limitations: lack of contextualization of classes (e.g., kitchen utensil instead of a bowl in Figure 16-left), incorrect questions or answers (e.g., about uncountable objects like grass in Figure 16-right), and classes in the WordNet hierarchy that are not overly specific (e.g., derby horse race) or generic (e.g., entity). Thus, while every image in VG enables many taxonomic questions to be created, it remains challenging to automatically distinguish high- from low-quality question-answer pairs. More fundamentally, creating question-answer pairs for other object properties (e.g., functional) and reasoning types (e.g., counterfactual) is hindered by the absence of annotation. While LLMs can be employed to generate questions for these categories, we established that these questions and answers are often incorrect or ambiguous (e.g., suggesting that buildings, doors, and windows can all provide shelter). In conclusion, we find that densely annotated image datasets remain insufficient to generate high-quality large-scale versions of OPTICS. Yet, they could potentially be used to speed up human annotation by providing an initial set of samples or for distant supervision of VQA models.